# Lipschitz-free spaces and Bossard’s reduction argument

### v1

@@ 3. The main results
Next, while not strictly necessary for what is to come, we determine the descriptive complexity of the classes of countably infinite complete discrete metric spaces and complete separable p1u metric spaces.
@@
which are closed in $\mathcal{R}\times\omega^{\omega}$ and $\omega^{\omega}$, respectively. Then $(d,(x_{i}))\in A:=\bigcap_{k\in\omega}\bigcup_{m\in\omega}A_{k,m}$ if and only if $(x_{i})\in\omega^{\omega}$ is Cauchy in $(\omega,d)$, and $(x_{i})\in B:=\bigcup_{m\in\omega}B_{m}$ if and only if $(x_{i})$ is eventually constant. Therefore $\mathcal{R}\setminus\mathcal{R}_{\mathrm{dis}}$ is $\mathbf{\Sigma}^{1}_{1}$ because it is the image of $A\setminus(\mathcal{R}\times B)\subset\mathcal{R}\times\omega^{\omega}$ under the projection onto the first coordinate. ∎
@@
Proposition 3.4 can be proved more directly using Kirchheim’s characterisation of p1u spaces mentioned above. Later, in Corollary 3.9, we upgrade $\mathcal{R}_{\mathrm{dis}}$ and $\mathcal{R}_{\mathrm{p1u}}$ to $\mathbf{\Pi}^{1}_{1}$-complete. For this and the main results we will need the following lemma, which is a tool for generating continuous maps from power sets to $\mathcal{R}$.
@@
Let $(M,d)$ be a metric space, $A\subset\omega$, ${\phi_{0}}:{A}\to{M}$ and ${\phi_{1},\phi_{2}}:{\omega\setminus A}\to{M}$ injections with pairwise disjoint ranges, and let ${\tau}:{\omega\setminus A}\to{\Gamma}$, where $\Gamma$ is a set. Then the map ${\Psi}:{2^{\Gamma}}\to{M^{\omega}}$ given by
@@
To check continuity of $\Psi$, let $E\in 2^{\Gamma}$ and let $F\subset\omega$ be finite. Define the finite sets $G=\tau(F\setminus A)\cap E$ and $H=\tau(F\setminus A)\setminus E$; then $G\subset E$, $H\cap E=\varnothing$ and $\Psi(E^{\prime})(n)=\Psi(E)(n)$ whenever $n\in F$, $E^{\prime}\in 2^{\Gamma}$, $G\subset E^{\prime}$ and $H\cap E^{\prime}=\varnothing$. Therefore $\Psi$ is continuous. As $\Delta$ is the composition of $\Psi$ and another continuous map, it is also continuous. Since $\phi_{0},\phi_{1},\phi_{2}$ are injective and have pairwise disjoint ranges, $\Psi(E)$ is injective for all $E\in 2^{\Gamma}$. It follows that $\Delta(E)\in\mathcal{R}$. The last assertion is obvious. ∎
@@
Given $\varepsilon>0$ and an infinite branch $B\subset\omega^{<\omega}$, there exists $n\in\omega$ such that
@@
Let $(M,\theta)$ be a separable metric space, let $(x_{i})$ be a dense sequence in $M$ whose elements repeat infinitely often, and let $\rho$ be a metric on $\omega^{<\omega}$ satisfying (3.1). Then the metric $d$ on $\omega^{<\omega}$ defined by
$d(s,t)=\rho(s,t)+\theta(x_{|s|},x_{|t|}),$
has the property that $(M,\theta)$ embeds isometrically into the completion $\overline{{(B,d)}}$ of $(B,d)$ whenever $B\subset\omega^{<\omega}$ is an infinite branch.
@@
Let $B\subset\omega^{<\omega}$ be an infinite branch. Then there exists $q\in\omega^{\omega}$ such that $B=\left\{{q\mathord{\upharpoonright}_{n}}\;:\;\,{n\in\omega}\right\}$. We define an isometric embedding ${\phi}:{M}\to{\overline{{(B,d)}}}$ as follows. Given $x\in M$, because the elements of $(x_{i})$ repeat infinitely often, there exist $i_{0}<i_{1}<i_{2}<\ldots$ such that $\theta(x_{i_{k}},x)\to 0$. Given that $\rho$ satisfies (3.1), $(q\mathord{\upharpoonright}_{i_{k}})$ is $\rho$-Cauchy, hence $(q\mathord{\upharpoonright}_{i_{k}})$ is also $d$-Cauchy. Let $\phi(x)=\lim_{k}q\mathord{\upharpoonright}_{i_{k}}\in\overline{{(B,d)}}$. This is well-defined because if $j_{0}<j_{1}<j_{2}<\ldots$ also satisfy $\theta(x_{j_{k}},x)\to 0$ then $d(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})=\rho(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})+\theta(x_{i_{k}},x_{j_{k}})\to 0$, again by (3.1). Finally, we check that $\phi$ is an isometric embedding. With $x$ and the $i_{k}$ as above, let $y\in M$ and this time let $j_{0}<j_{1}<j_{2}<\ldots$ such that $\theta(y,x_{j_{k}})\to 0$. Then once more by (3.1),
$d(\phi(x),\phi(y))=\lim_{k}d(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})=\lim_{k}\rho(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})+\theta(x_{i_{k}},x_{j_{k}})=\theta(x,y).\qed$
@@
Let $d$ be the metric from Lemma 3.7, where $(M,\theta)$ is arbitrary and $\rho$ is as above. Then $(\omega^{<\omega},d)$ is discrete, and $(T,d)$ is in addition complete whenever $T\in\mathbf{WF}$.
@@
$d(s,t)\geqslant\rho(s,t)=2^{-|s\wedge t|-1}\geqslant 2^{-|s|-1}.$
@@
Hence $d(s,t)\geqslant 2^{-|s|-1}$ whenever $t\in\omega^{<\omega}\setminus\{s\}$, meaning that $d$ is discrete.
@@
Now let $T\in\mathbf{WF}$ and let $(s_{n})\subset T$ be $d$-Cauchy. We find a strictly increasing finite sequence $r_{0}\prec r_{1}\prec r_{2}\prec\ldots$ of elements of $\omega^{<\omega}$ and infinite subsets $L_{0}\supset L_{1}\supset L_{2}\supset\ldots$ of $\omega$ such that $r_{k}\preccurlyeq s_{n}$ for all $n\in L_{k}$. Begin by setting $r_{0}=\varnothing$ and $L_{0}=\omega$. Given $r_{k}$ and $L_{k}$, either there exists $m\in\omega$ and an infinite subset $L_{k+1}\subset L_{k}$ such that $r_{k}^{\frown}m\preccurlyeq s_{n}$ for all $n\in L_{k+1}$, or not. If the first case holds, set $r_{k+1}=r_{k}^{\frown}m$ and continue the process, else terminate it. Because $T\in\mathbf{WF}$, this process must terminate at some finite stage $k$. By construction we know that $L_{k}$ is infinite and $r_{k}\preccurlyeq s_{n}$ for all $n\in L_{k}$ but also, given $m\in\omega$, the set $H_{m}:=\left\{{n\in L_{k}}\;:\;\,{r_{k}^{\frown}m\preccurlyeq s_{n}}\right\}$ must be finite. If $n_{1}\in H_{m_{1}}$ and $n_{2}\in H_{m_{2}}$ for distinct $m_{1},m_{2}\in\omega$, then $s_{n_{1}}\perp s_{n_{2}}$ and $s_{n_{1}}\wedge s_{n_{2}}=r_{k}$, giving
$d(s_{n_{1}},s_{n_{2}})\geqslant\rho(s_{n_{1}},s_{n_{2}})=2^{-|r_{k}|-1}.$
@@
Because $(s_{n})$ is $d$-Cauchy, it follows that $H_{m}$ is empty for all but finitely many $m\in\omega$, and thus $s_{n}=r_{k}$ for infinitely many $n\in L_{k}$. Again because $(s_{n})$ is $d$-Cauchy, we conclude that it must converge to $r_{k}$. This proves that $(T,d)$ is complete. ∎
The proofs of Theorem 1.1 and Theorem 1.2 use Bossard’s reduction argument [bossard:02]*Theorem 3.2.
@@
Let $X$ be a separable Banach space. By Proposition 3.2, $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. Assume that $\mathcal{R}_{\mathrm{dis}}\subset\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$. Let $\theta$ be the usual metric on $B_{c_{0}}$ inherited from $c_{0}$, and let $(\omega^{<\omega},d)$ be the metric space obtained from Lemma 3.7 applied to $(B_{c_{0}},\theta)$. It is evident that $d\leqslant\frac{5}{2}$. Let $M=\omega^{<\omega}\cup(\omega^{<\omega}\times\{0\})$ and extend $d$ to a metric on $M$ by setting $d(x,y)=\frac{5}{2}$ whenever $x\neq y$ and $\{x,y\}\setminus\omega^{<\omega}$ is non-empty (it is straightforward to check that the extended function satisfies the triangle inequality).
Set $A=\phi_{0}=\varnothing$, $\Gamma=\omega^{<\omega}$, let ${\tau}:{\omega}\to{\Gamma}$ be a bijection, and define injections ${\phi_{1},\phi_{2}}:{\omega}\to{M}$ by $\phi_{1}(n)=(\tau(n),0)$ and $\phi_{2}(n)=\tau(n)$, $n\in\omega$. Let ${\Delta}:{2^{\Gamma}}\to{\mathcal{R}}$ be the continuous map furnished by Lemma 3.5, which we then restrict to $\mathbf{Tr}$.
Let $T\in\mathbf{Tr}$. We observe that $\Psi(T)(\omega)=T\cup((\omega^{<\omega}\setminus T)\times\{0\})$. Using Lemma 3.8, $(T,d)$ is discrete and complete if $T\in\mathbf{WF}$. Hence $(\omega,\Delta(T))$, which is isometric to $(\Psi(T)(\omega),d)$, is also discrete and complete because every point in $(\omega^{<\omega}\setminus T)\times\{0\}$ is $d$-separated from everything else by $\frac{5}{2}$. Therefore $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$. Now let $T\in\mathbf{IF}$ and let $B\subset T$ be an infinite branch. Then the completion $M_{\Delta(T)}$ of $(\omega,\Delta(T))$ contains an isometric copy of the completion of $(B,d)$. By Lemma 3.7, $B_{c_{0}}\hookrightarrow M_{\Delta(T)}$.
As $\mathcal{R}_{\mathrm{dis}}\subset\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$, we have $\mathbf{WF}=\Delta^{-1}(\mathcal{R}_{\mathrm{dis}})\subset\Delta^{-1}(\langle X\rangle^{\mathcal{R}}_{\hookrightarrow})$. Since $\mathbf{WF}$ is $\mathbf{\Pi}^{1}_{1}$-complete and $\Delta^{-1}(\langle X\rangle^{\mathcal{R}}_{\hookrightarrow})$ is $\mathbf{\Sigma}^{1}_{1}$, there exists $T\in\mathbf{IF}$ such that $\Delta(T)\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$. Therefore $\mathcal{F}({B_{c_{0}}})\hookrightarrow\mathcal{F}({M_{d}})\hookrightarrow X$. Because $Y\hookrightarrow\mathcal{F}({B_{c_{0}}})$ whenever $Y$ is a separable Banach space ([dutrieux:ferenczi:06]*p. 1042 and [kaufmann:15]*Corollary 3.3), the conclusion follows. ∎

### v2

@@ 3. Three main results
Next, we determine the descriptive complexity of the classes of countably infinite complete discrete metric spaces and complete separable p1u metric spaces.
@@
which are closed in $\mathcal{R}\times\omega^{\omega}$ and $\omega^{\omega}$, respectively. Then $(d,(x_{i}))\in A:=\bigcap_{k\in\omega}\bigcup_{m\in\omega}A_{k,m}$ if and only if $(x_{i})\in\omega^{\omega}$ is Cauchy in $(\omega,d)$, and $(x_{i})\in B:=\bigcup_{m\in\omega}B_{m}$ if and only if $(x_{i})$ is eventually constant. Therefore $\mathcal{R}\setminus\mathcal{R}_{\mathrm{dis}}$ is $\mathbf{\Sigma}^{1}_{1}$ because it is the image of $A\setminus(\mathcal{R}\times B)\subseteq\mathcal{R}\times\omega^{\omega}$ under the projection onto the first coordinate. ∎
@@
Proposition 3.4 can be proved more directly using Kirchheim’s characterisation of p1u spaces mentioned above. Later, in Proposition 3.13, we upgrade $\mathcal{R}_{\mathrm{dis}}$ and $\mathcal{R}_{\mathrm{p1u}}$ to $\mathbf{\Pi}^{1}_{1}$-complete. For this and the main results we will need the following lemma, which is a tool for generating continuous maps from power sets to $\mathcal{R}$.
@@
Let $(M,d)$ be a metric space, $A\subseteq\omega$, ${\phi_{0}}:{A}\to{M}$ and ${\phi_{1},\phi_{2}}:{\omega\setminus A}\to{M}$ injections with pairwise disjoint ranges, and let ${\tau}:{\omega\setminus A}\to{\Gamma}$, where $\Gamma$ is a set. Then the map ${\Psi}:{2^{\Gamma}}\to{M^{\omega}}$ given by
@@
To check continuity of $\Psi$, let $E\in 2^{\Gamma}$ and let $F\subseteq\omega$ be finite. Define the finite sets $G=\tau(F\setminus A)\cap E$ and $H=\tau(F\setminus A)\setminus E$; then $G\subseteq E$, $H\cap E=\varnothing$ and $\Psi(E^{\prime})(n)=\Psi(E)(n)$ whenever $n\in F$, $E^{\prime}\in 2^{\Gamma}$, $G\subseteq E^{\prime}$ and $H\cap E^{\prime}=\varnothing$. Therefore $\Psi$ is continuous. As $\Delta$ is the composition of $\Psi$ and another continuous map, it is also continuous. Since $\phi_{0},\phi_{1},\phi_{2}$ are injective and have pairwise disjoint ranges, $\Psi(E)$ is injective for all $E\in 2^{\Gamma}$. It follows that $\Delta(E)\in\mathcal{R}$. The last assertion is obvious. ∎
@@
Given $\varepsilon>0$ and an infinite branch $B\subseteq\omega^{<\omega}$, there exists $n\in\omega$ such that
@@
In the next result and a number of others, $\overline{{M}}$ shall denote the completion of the metric space $M$.
@@
Let $M=(M,\theta)$ be a separable metric space, let $(x_{i})_{i\in\omega}$ be a dense sequence in $M$ whose elements repeat infinitely often, let $\rho$ be a metric on $\omega^{<\omega}$ satisfying (3.1) and define a second metric $\eta$ on $\omega^{<\omega}$ by
$\eta(s,t)=\rho(s,t)+\theta(x_{|s|},x_{|t|}).$
Then whenever $T\in\mathbf{IF}$ and $B\subseteq T$ is an infinite branch, there exists an isometric embedding ${\phi}:{M}\to{\overline{{(B,\eta)}}\subseteq\overline{{(T,\eta)}}}$. Moreover, if $M$ is complete then there is a 1-Lipschitz retraction ${r}:{\overline{{(T,\eta)}}}\to{\phi(M)}$.
@@
Let $q\in\omega^{\omega}$ such that $B=\left\{{q\mathord{\upharpoonright}_{n}}\;:\;\,{n\in\omega}\right\}$. We define an isometric embedding ${\phi}:{M}\to{\overline{{(B,\eta)}}}$ as follows. Given $x\in M$, because the elements of $(x_{i})$ repeat infinitely often, there exist $i_{0}<i_{1}<i_{2}<\ldots$ such that $\theta(x_{i_{k}},x)\to 0$. Given that $\rho$ satisfies (3.1), $(q\mathord{\upharpoonright}_{i_{k}})$ is $\rho$-Cauchy, hence $(q\mathord{\upharpoonright}_{i_{k}})$ is also $\eta$-Cauchy. Let $\phi(x)=\lim_{k}q\mathord{\upharpoonright}_{i_{k}}\in\overline{{(B,\eta)}}$. This is well-defined because if $j_{0}<j_{1}<j_{2}<\ldots$ also satisfy $\theta(x_{j_{k}},x)\to 0$ then $\eta(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})=\rho(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})+\theta(x_{i_{k}},x_{j_{k}})\to 0$, again by (3.1). Next, we check that $\phi$ is an isometric embedding. With $x$ and the $i_{k}$ as above, let $y\in M$ and this time let $j_{0}<j_{1}<j_{2}<\ldots$ such that $\theta(y,x_{j_{k}})\to 0$. Then once more by (3.1),
$\eta(\phi(x),\phi(y))=\lim_{k}\eta(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})=\lim_{k}\rho(q\mathord{\upharpoonright}_{i_{k}},q\mathord{\upharpoonright}_{j_{k}})+\theta(x_{i_{k}},x_{j_{k}})=\theta(x,y).$
Finally, assume $M$ is complete. Define ${\zeta}:{T}\to{M}$ by $\zeta(s)=x_{|s|}$. This map is $1$-Lipschitz because $\theta(\zeta(s),\zeta(t))=\theta(x_{|s|},x_{|t|})\leqslant\eta(s,t)$, $s,t\in T$. Let ${r}:{\overline{{(T,\eta)}}}\to{\overline{{(B,\eta)}}}$ be the $1$-Lipschitz extension of $\phi\circ\zeta$ to $\overline{{(T,\eta)}}$. By the continuity of $r$ and the completeness of $M$, we have $r(\overline{{(T,\eta)}})\subseteq\phi(M)$. Moreover $r(\phi(x))=\phi(x)$ whenever $x\in M$. Indeed, given $i_{0}<i_{1}<i_{2}<\ldots$ such that $\theta(x_{i_{k}},x)\to 0$, we have $\phi(x)=\lim_{k}q\mathord{\upharpoonright}_{i_{k}}$, so $r(\phi(x))=\lim_{k}\phi(\zeta(q\mathord{\upharpoonright}_{i_{k}}))=\lim_{k}\phi(x_{i_{k}})=\phi(x)$. ∎
@@
Let $\eta$ be the metric from Lemma 3.7, where $(M,\theta)$ is arbitrary and $\rho$ is as above. Then $(\omega^{<\omega},\eta)$ is discrete, and in addition $(T,\eta)$ is complete whenever $T\in\mathbf{WF}$.
@@
$\eta(s,t)\geqslant\rho(s,t)=2^{-|s\wedge t|-1}\geqslant 2^{-|s|-1}.$
@@
Hence $\eta(s,t)\geqslant 2^{-|s|-1}$ whenever $t\in\omega^{<\omega}\setminus\{s\}$, meaning that $\eta$ is discrete.
@@
Now let $T\in\mathbf{WF}$ and let $(s_{n})\subseteq T$ be $\eta$-Cauchy. We find a strictly increasing finite sequence $r_{0}\prec r_{1}\prec r_{2}\prec\ldots$ of elements of $\omega^{<\omega}$ and infinite subsets $L_{0}\supseteq L_{1}\supseteq L_{2}\supseteq\ldots$ of $\omega$ such that $r_{k}\preccurlyeq s_{n}$ for all $n\in L_{k}$. Begin by setting $r_{0}=\varnothing$ and $L_{0}=\omega$. Given $r_{k}$ and $L_{k}$, either there exists $m\in\omega$ and an infinite subset $L_{k+1}\subseteq L_{k}$ such that $r_{k}^{\frown}m\preccurlyeq s_{n}$ for all $n\in L_{k+1}$, or not. If the first case holds, set $r_{k+1}=r_{k}^{\frown}m$ and continue the process, else terminate it. Because $T\in\mathbf{WF}$, this process must terminate at some finite stage $k$. By construction we know that $L_{k}$ is infinite and $r_{k}\preccurlyeq s_{n}$ for all $n\in L_{k}$ but also, given $m\in\omega$, the set $H_{m}:=\left\{{n\in L_{k}}\;:\;\,{r_{k}^{\frown}m\preccurlyeq s_{n}}\right\}$ must be finite. If $n_{1}\in H_{m_{1}}$ and $n_{2}\in H_{m_{2}}$ for distinct $m_{1},m_{2}\in\omega$, then $s_{n_{1}}\perp s_{n_{2}}$ and $s_{n_{1}}\wedge s_{n_{2}}=r_{k}$, giving
$\eta(s_{n_{1}},s_{n_{2}})\geqslant\rho(s_{n_{1}},s_{n_{2}})=2^{-|r_{k}|-1}.$
@@
Because $(s_{n})$ is $\eta$-Cauchy, it follows that $H_{m}$ is empty for all but finitely many $m\in\omega$, and thus $s_{n}=r_{k}$ for infinitely many $n\in L_{k}$. Again because $(s_{n})$ is $\eta$-Cauchy, we conclude that it must converge to $r_{k}$. This proves that $(T,\eta)$ is complete. ∎
We make use of the last three lemmas in the next result. Lemmas 3.9 and 3.10 below are our main tools.
Lemma 3.9.
Let $M=(M,\theta)$ be a separable metric space. There exists a continuous map ${\Delta}:{\mathbf{Tr}}\to{\mathcal{R}}$ such that $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$ whenever $T\in\mathbf{WF}$ and $M$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$ whenever $T\in\mathbf{IF}$.
Proof.
Let $(\omega^{<\omega},\eta)$ be the metric space obtained from Lemma 3.7 applied to $M$. Let $W=\omega^{<\omega}\cup(\omega^{<\omega}\cup\{0\})$. We extend $\eta$ to a metric on $W$, again labelled $\eta$, by setting
$\eta(x,y)=\begin{cases}0&\text{if }x,y\notin\omega^{<\omega},\,x=y,\\ 1&\text{if }x,y\notin\omega^{<\omega},\,x\neq y,\\ 1+\eta(\varnothing,x)&\text{if }x\in\omega^{<\omega},\,y\notin\omega^{<\omega},\\ 1+\eta(\varnothing,y)&\text{if }x\notin\omega^{<\omega},\,y\in\omega^{<\omega}.\end{cases}$
It is routine to verify that this extension satisfies the triangle inequality.
Set $A=\phi_{0}=\varnothing$, $\Gamma=\omega^{<\omega}$, let ${\tau}:{\omega}\to{\Gamma}$ be a bijection, and define injections ${\phi_{1},\phi_{2}}:{\omega}\to{W}$ by $\phi_{1}(n)=(\tau(n),0)$ and $\phi_{2}(n)=\tau(n)$, $n\in\omega$. Let ${\Delta}:{2^{\Gamma}}\to{\mathcal{R}}$ be the continuous map furnished by Lemma 3.5, which we then restrict to $\mathbf{Tr}$. We observe that $\Psi(T)(\omega)=T\cup((\omega^{<\omega}\setminus T)\times\{0\})$ for all $T\in\mathbf{Tr}$.
Let $T\in\mathbf{WF}$. Using Lemma 3.8, $(T,\eta)$ is discrete and complete. Hence $(\omega,\Delta(T))$, which is isometric to $(\Psi(T)(\omega),\eta)$, is also discrete and complete because every point in $(\omega^{<\omega}\setminus T)\times\{0\}$ is $\eta$-separated from everything else by at least $1$. Therefore $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$.
Now let $T\in\mathbf{IF}$. We claim that $M$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$. First of all, the function ${r^{\prime}}:{\Psi(T)(\omega)}\to{T}$, defined by
$r^{\prime}(x)=\begin{cases}x&\text{if }x\in T,\\ \varnothing&\text{if }x\notin T,\end{cases}$
is a $1$-Lipschitz retraction of $(\Psi(T)(\omega),\eta)$ onto $(T,\eta)$. We extend this to a $1$-Lipschitz retraction ${r^{\prime}}:{\overline{{(\Psi(T)(\omega),\eta)}}}\to{\overline{{(T,\eta)}}}$. According to Lemma 3.7, there is an isometric embedding ${\phi}:{M}\to{\overline{{(T,\eta)}}}$ and a $1$-Lipschitz retraction ${r}:{\overline{{(T,\eta)}}}\to{\phi(M)}$. Therefore ${r\circ r^{\prime}}:{\overline{{(\Psi(T)(\omega),\eta)}}}\to{\phi(M)}$ is a 1-Lipschitz retraction. Since $M_{\Delta(T)}\equiv\overline{{(\Psi(T)(\omega),\eta)}}$, our claim is proved. ∎
The final lemma below uses Bossard’s reduction argument [bossard:02]*Theorem 3.2; Theorems 1.11.3 emerge as relatively straightforward consequences of it.
Lemma 3.10.
Let $\mathcal{A}\subseteq\mathcal{R}$ have the property that $d^{\prime}\in\mathcal{A}$ whenever $d^{\prime}\in\mathcal{R}$ and $M_{d^{\prime}}$ is isometric to a $1$-Lipschitz retract of $M_{d}$ for some $d\in\mathcal{A}$. Assume that $\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A}$ is $\mathbf{\Sigma}^{1}_{1}$. Then $\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A}=\mathcal{R}$.
Proof.
Let $d\in\mathcal{R}$. Apply Lemma 3.9 to $M:=M_{d}$. The assumption $\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A}$ is $\mathbf{\Sigma}^{1}_{1}$ implies $\Delta^{-1}(\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A})$ is also $\mathbf{\Sigma}^{1}_{1}$. As $\mathbf{WF}\subseteq\Delta^{-1}(\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A})$ and $\mathbf{WF}$ is $\mathbf{\Pi}^{1}_{1}$-complete, by Suslin’s theorem there exists $T\in\mathbf{IF}$ such that $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A}$. Because $T\in\mathbf{IF}$, $M_{d}$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$. Given our assumption about $\mathcal{A}$ and the fact that $\mathcal{R}_{\mathrm{dis}}$ also enjoys this property, it follows that $d\in\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A}$. Therefore $\mathcal{R}_{\mathrm{dis}}\cup\mathcal{A}=\mathcal{R}$. ∎
@@
Let $X$ be a separable Banach space. By Proposition 3.2, $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. Let $d^{\prime}\in\mathcal{R}$ and $d\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$ such that $M_{d^{\prime}}$ is isometric to a 1-Lipschitz retract of $M_{d}$. By [weaver:18]*Theorems 3.6 and 3.7, $\mathcal{F}({M_{d^{\prime}}})\hookrightarrow\mathcal{F}({M_{d}})$ and thus $d^{\prime}\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$.
Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$. By Lemma 3.10, $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}=\mathcal{R}$. As $Y\hookrightarrow\mathcal{F}({c_{0}})$ whenever $Y$ is a separable Banach space ([dutrieux:ferenczi:06]*p. 1042) and $c_{0}\equiv M_{d}$ for some $d\in\mathcal{R}$, it follows that $Y\hookrightarrow X$ for all such $Y$. ∎